# Artificial Inductive Bias for Synthetic Tabular Data Generation in Data-Scarce Scenarios

### v1

@@ 2.1 Transfer learning
Transfer learning is a machine learning paradigm that leverages knowledge acquired from a context domain (also called the source domain) to enhance learning performance in a new target domain [30]. This approach aims to improve the learning process in the target domain by capitalizing on the knowledge gained from solving related tasks in the context domain. This technique has demonstrated its efficacy in fields where data scarcity is a common challenge, such as the medical field [31].
@@
Formally, based on the definition in [30], we can define a domain $\mathcal{D}$ by a feature space $\mathcal{X}$ and a marginal probability distribution $p(x_{r})$. Two domains are considered distinct if their feature spaces $\mathcal{X}_{1},\mathcal{X}_{2}$ or marginal probability distributions $p(x_{1}),p(x_{2})$ differ, i.e., if $\mathcal{X}_{1}\neq\mathcal{X}_{2}$ or $p(x_{1})\neq p(x_{2})$. The core objective of transfer learning is to leverage the knowledge learned in a context domain $\mathcal{D}_{context}$ to improve learning in a target domain $\mathcal{D}_{target}$. This is typically achieved when the context and target domains differ, i.e., $\mathcal{D}_{context}\neq\mathcal{D}_{target}$.
@@
Our work focuses on a scenario where the context domain $\mathcal{D}_{context}$ consists of data $x_{g}$ generated by a DGM. The target domain $\mathcal{D}_{target}$, on the other hand, consists of $x_{r}$. Our approach leverages the representational power learned by the DGM $p_{\theta}$ on $x_{g}$ to provide a strong starting point for learning in the target domain with real data $x_{r}$. This transfer of knowledge is achieved by initializing the model weights for the target domain with the weights learned from the DGM model trained on the generated data.
@@
Transfer learning can be categorized into homogeneous and heterogeneous settings based on the feature spaces of the domains [32]. Homogeneous transfer learning applies when the context and target domains share the same feature space $\mathcal{X}_{1}=\mathcal{X}_{2}$, while heterogeneous transfer learning deals with scenarios where feature spaces differ $\mathcal{X}_{1}\neq\mathcal{X}_{2}$. This work focuses on homogeneous transfer learning, where the context domain serves as an augmented version of the target domain. The key difference between the domains in our case lies in the number of samples, leading to situations where the empirical distributions of the data differ, i.e., $p_{\theta}(x_{g})\neq p(x_{r})$.
@@
Within homogeneous transfer learning, various methodologies exist to improve target task performance by capitalizing on knowledge from a related source domain. These techniques encompass instance-based [33], relational knowledge transfer [34], feature-based [35], and, as employed in this work, parameter-based [36] transfer through shared model parameters or hyperparameter distributions. This study leverages a two-stage parameter-based transfer learning approach. The first stage involves either pre-training or model averaging, followed by fine-tuning in the second stage. Subsequent sections will delve deeper into both pre-training and model averaging techniques. Upon completion of one of these initial phases, fine-tuning serves to refine the model parameters, ultimately achieving optimal adaptation for the target domain.
@@ 2.1.1 Pre-training
Figure 2: Block diagram for the pre-training case. The inductive bias is introduced by training a DGM $p_{\theta_{pt}}$ on a large collection of $x_{g}$ samples. The weights learned from this training process with abundant samples serve as the initial parameters $\hat{\theta}$ for the fine-tunning process using the real data $x_{r}$ to obtain $p_{\hat{\theta}}$.
Pre-training is a frequently adopted strategy for introducing an inductive bias into a model. By leveraging a pre-trained model on a context domain, the target model gains generalizable features that enhance its performance on a target domain. However, while pre-training is a standard in computer vision and natural language processing, achieving similar success with tabular data remains a challenge. This disparity arises from the inherent heterogeneity of the features of the tables, which creates substantial feature space shifts between pre-training and downstream datasets, hindering effective knowledge transfer. Despite these challenges, recent efforts like [37] and [38] explore tabular transfer learning with promising results. Although these studies demonstrate potential, achieving comprehensive parameter transfer in tabular data requires further research to establish best practices and unlock the full potential of pre-training in this domain.
In this work, pre-training involves the following steps. First, we train a separate DGM $p_{\theta_{pt}}$ using synthetic data $x_{g}$ as training data. Since $x_{g}$ is sampled from the initial DGM, $x_{g}\sim p_{\theta}$, we can generate a vast amount of synthetic data. This abundance circumvents the limitations associated with training in small datasets, such as overfitting. Then, the optimal weights $\theta_{pt}^{*}$ from DGM $p_{\theta_{pt}}$ are used as initial weights $\theta_{0}$ to train the generative model $p_{\hat{\theta}}$ (see Figure 1).
In essence, our approach aligns with the well-established concept of data augmentation. We generate synthetic data $x_{g}$, which may not perfectly capture the intricacies of the original data $x_{r}$. However, we used these synthetic data to train another DGM $p_{\theta_{pt}}$. Although $p_{\theta_{pt}}$ might generate lower quality synthetic samples, our objective is to exploit the information encoded within this DGM to establish an initial set of weights for the DGM that will eventually be trained on $x_{r}$. In other words, we exploit the knowledge of the generative model $p_{\theta_{pt}}$, the context domain, to obtain a better generative model $p_{\hat{\theta}}$, which is our target domain. Figure 2 provides a visual representation of this pre-training procedure.
@@ 2.1.2 Model Averaging
Figure 3: Block diagram for the model averaging case. The inductive bias is introduced by training a DGM $p_{\theta}$ on $x_{r}$ using $S$ different seeds. The average of the weights learned from these training processes serve as the initial parameters $\hat{\theta}$ for the fine-tunning process using the real data $x_{r}$ to obtain $p_{\hat{\theta}}$.
The concept of model averaging emerged in the 1960s, primarily within the field of economics [39, 40]. Traditional empirical research often selects a single “best" model after searching a wide space of possibilities. However, this approach can underestimate the real uncertainty and lead to overly confident conclusions. Model averaging offers a compelling alternative. By combining multiple models, the resulting ensemble can outperform any individual model. This approach aligns with the core principles of statistical modeling: maximizing information use and balancing flexibility with overfitting. In essence, model averaging extends the concept of model selection by leveraging insights from all the models considered.
While pre-training can be incorporated with any DGM, our approach focuses on models where the training process is sensitive to initial conditions, such as VAEs. In such cases, it is common to train the DGM $p_{\theta}$ with multiple initial conditions (seeds) and potentially discard “bad" seeds based on a specific metric. We propose using these discarded seeds to create an artificial inductive bias. The simplest implementation involves averaging the model parameters. In this case, our context domains are the different results of each seed, and the target domain is obtained by averaging across the context domains. If we train $S$ different seeds for $p_{\theta}$, resulting in $S$ models with parameters $\theta_{s}$, we propose using the average of these weights as the inductive bias:
@@
This straightforward approach is computationally efficient, requiring only the calculation of the average across the precomputed weights. It assumes that the average model may capture a robust inductive bias, leading to improved performance. Figure 3 summarizes this process.
@@ 2.2 Meta-learning
Traditional machine learning models often rely on large volumes of data to achieve optimal performance in specific tasks. In contrast, meta-learning introduces a distinct paradigm by training algorithms with the ability to “learn to learn" [41], enabling them to rapidly adapt to new tasks with minimal data. This departure from the conventional requirement of extensive datasets for each new task allows meta-learning algorithms to leverage knowledge gained from addressing numerous related tasks. Through introspective analysis of past experiences, these models dynamically adjust their learning strategies when confronted with novel situations, making them more efficient learners and requiring less data to perform well on tasks with similar characteristics.
In this work, we exploit the multi-seed training configuration of certain DGMs. By treating each of the $S$ different seeds obtained after training the DGM as a distinct task, we construct a meta-learning framework.
@@ 2.2.1 MAML
MAML is a prevalent approach within the field of meta-learning [42]. It identifies the initial set of weights denoted by $\theta_{MAML}$ by leveraging various tasks, enabling rapid and data-efficient adaptation to new tasks. This efficiency comes from the ability to fine-tune $\theta_{MAML}$ with minimal data for each new task. However, successful application of MAML requires access to a diverse collection of tasks for effective learning.
Formally, we can frame the problem by starting with a common single-task learning scenario and transforming it into the meta-learning framework. Consider a task $\mathcal{T}$ that consists of an input $x$ sampled from a probability distribution $\mathcal{D}$. For simplicity, we define a task instance $\mathcal{T}$ as a tuple comprising a dataset $\mathcal{D}$ and its corresponding loss function $\mathcal{L}$. To solve the task $\mathcal{T}$ we need to obtain an optimal model parameterized by a task-specific parameter $\omega^{*}$, which minimizes a loss function $\mathcal{L}$ on the data of the task as follows:
In single-task learning, hyperparameter optimization is achieved by splitting the dataset $\mathcal{D}$ into two disjoint subsets $\mathcal{D}=\mathcal{D}^{(t)}\cup\mathcal{D}^{(v)}$, which are the training and validation sets, respectively. The meta-learning setting aims to develop a general-purpose learning algorithm that excels across a distribution of tasks represented by $p(\mathcal{T})$ [43]. The objective is to use training tasks to train a meta-learning model $\theta_{MAML}$ that can be fine-tuned to obtain $\omega$ to perform well on unseen tasks sampled from the same task environment $p(\mathcal{T})$. Meta-learning methods utilize meta-parameters to model the common latent structure of the task distribution $p(\mathcal{T})$. Therefore, we consider meta-learning as an extension of hyperparameter optimization, where the hyperparameter of interest – often called a meta-parameter – is shared across many tasks.
In this work, the distribution of tasks is defined by the set of $S$ training seeds obtained after training the DGM. Given a set of $S$ training seeds following $p(\mathcal{T})$, each task $\mathcal{T}\sim p(\mathcal{T})$ is therefore formalized as $\mathcal{T}=\{\mathcal{D},\mathcal{L}\}$. Here, each dataset $\mathcal{D}$ consists of synthetic data points $x_{g}^{s}$ drawn from the model for the different training seeds. The loss function $\mathcal{L}$ corresponds to the DGM loss function. The specific form of $\mathcal{L}$ depends on the chosen DGM. If the chosen DGM is a VAE the loss function $\mathcal{L}$ would be the negative of the Evidence Lower BOund (ELBO) [44]. In contrast, if a GAN is used, the loss function $\mathcal{L}$ would be the minimax loss function arising from the interplay between the generator and discriminator networks [45]. It is important to note that both VAEs and GANs use two neural networks within their architecture, different from the single network architectures commonly found in state-of-the-art applications [46, 47].
Solving this problem using the MAML approach requires access to a collection of $B$ tasks sampled from $p(\mathcal{T})$. We denote this set of tasks $\mathcal{T}_{b}$ used for training as $\mathcal{D}_{b}=\{(\mathcal{D}_{b}^{(t)},\mathcal{D}^{(v)}_{b})\}_{b=1}^{B}$, where each task $b$ has dedicated meta-training and meta-validation data, respectively. The goal of meta-training is to find the optimal $\omega^{*}_{b}$ for a given each task $b$ given $\theta_{MAML}$. This $\theta_{MAML}$ essentially captures the ability to learn effectively from new data. In this context, the task-related parameter $\omega_{b}$ denotes the parameters of the two different networks that comprise the VAE, i.e. the encoder and decoder’s task-specific parameters. After meta-training, the learned $\omega^{*}_{b}$ is used to guide the training of a base model $\theta_{MAML}$. This procedure is called meta-testing. This essentially means that the model leverages the knowledge gained from previous tasks to improve the efficiency of learning on new tasks. This can be viewed as a bi-level optimization problem [48]:
@@
This equation essentially minimizes the expected loss across all tasks on the meta-validation sets, subject to the constraint that for each task, the task-specific parameter $\omega$ is optimized on the corresponding meta-training data.
Since in our work we are upgrading the parameters using gradient descent, we can reformulate Equation 4 as follows:
@@
Figure 4 illustrates the integration of the MAML procedure within the framework of our proposed methodology. In this context, the task space denoted by $p(\mathcal{T})$ corresponds to the various seeds $S$ obtained during the training process. Essentially, the task space encompasses the different probability distributions $p_{\theta_{s}}$ associated with each training seed. Ultimately, the meta-training steps lead to the identification of the desired parameters, denoted by $\theta_{MAML}$. Note that what $\theta_{MAML}$ represents is a set of parameters that adapt fast to new data; in our case, it means that the DGM initial parameters are chosen so that they adapt fast to generate real data.

### v2

@@ 2.1 Transfer learning
Transfer learning is a machine learning paradigm that leverages knowledge acquired from a context domain (also called the source domain) to enhance learning performance in a new target domain (Pan2010, ). This approach aims to improve the learning process in the target domain by leveraging knowledge gained from solving related tasks in the context domain. This technique has demonstrated its efficacy in fields where data scarcity is a common challenge, such as the medical field (Kim2022, ).
@@
Formally, based on the definition in Pan2010 , we can define a domain $\mathcal{D}$ by a feature space $\mathcal{X}$ and a marginal probability distribution $p(x_{r})$. Two domains are considered distinct if their feature spaces $\mathcal{X}_{1},\mathcal{X}_{2}$ or marginal probability distributions $p(x_{1}),p(x_{2})$ differ, i.e., if $\mathcal{X}_{1}\neq\mathcal{X}_{2}$ or $p(x_{1})\neq p(x_{2})$. The core objective of transfer learning is to leverage the knowledge learned in a context domain $\mathcal{D}_{context}$ to improve learning in a target domain $\mathcal{D}_{target}$. This is typically achieved when the context and target domains differ, i.e., $\mathcal{D}_{context}\neq\mathcal{D}_{target}$.
@@
Our work focuses on a scenario where the context domain $\mathcal{D}_{context}$ consists of data $x_{g}$ generated by a DGM. On the other hand, the target domain $\mathcal{D}_{target}$ consists of $x_{r}$. Our approach leverages the representational power learned by the DGM $p_{\theta}$ on $x_{g}$ to provide a strong starting point for learning in the target domain with real data $x_{r}$. This knowledge transfer is achieved by initializing the model weights for the target domain with the weights learned from the DGM model trained on the generated data.
@@
Transfer learning can be categorized into homogeneous and heterogeneous settings based on the feature spaces of the domains (weiss2016survey, ). Homogeneous transfer learning applies when the context and target domains share the same feature space $\mathcal{X}_{1}=\mathcal{X}_{2}$, while heterogeneous transfer learning deals with scenarios where feature spaces differ $\mathcal{X}_{1}\neq\mathcal{X}_{2}$. This work focuses on homogeneous transfer learning, where the context domain is an augmented version of the target domain. The key difference between the domains in our case lies in the number of samples, leading to situations where the empirical distributions of the data differ, i.e., $p_{\theta}(x_{g})\neq p(x_{r})$.
@@
Within homogeneous transfer learning, various methodologies exist to improve target task performance by leveraging knowledge from a related source domain. These techniques encompass instance-based (Chawla_2002, ), relational knowledge transfer (li-etal-2012-cross, ), feature-based (Long2014AdaptationRA, ), and, as employed in this work, parameter-based (Duan2012ExploitingWI, ) transfer through shared model parameters or hyperparameter distributions. This study leverages a two-stage parameter-based transfer learning approach. The first stage involves pre-training or model averaging, followed by fine-tuning in the second stage. Subsequent sections will delve deeper into both pre-training and model averaging techniques. Upon completing one of these initial phases, fine-tuning refines the model parameters, ultimately achieving optimal adaptation for the target domain.
@@ 2.1.1 Pre-training
Pre-training is a frequently adopted strategy for introducing an inductive bias into a model. By leveraging a pre-trained model on a context domain, the target model gains generalizable features that enhance its performance on a target domain. However, while pre-training is a standard in computer vision and natural language processing, achieving similar success with tabular data remains challenging. This disparity arises from the inherent heterogeneity of the features in the tables, which creates substantial shifts in feature space between pre-training and downstream datasets, hindering effective knowledge transfer. Despite these challenges, efforts such as wang2022transtab and zhu2023xtab have explored tabular transfer learning with promising results. Although these studies demonstrate potential, achieving comprehensive parameter transfer in tabular data requires further research to establish best practices and fully leverage the potential of pre-training in this domain.
In this work, pre-training involves the following steps. First, we train a separate DGM $p_{\theta_{pt}}$ using synthetic data $x_{g}$ as training data. Since $x_{g}$ is sampled from the initial DGM, $x_{g}\sim p_{\theta}$, we can generate a vast amount of synthetic data. This abundance circumvents the limitations associated with training in small datasets, such as overfitting. Then, the optimal weights $\theta_{pt}^{*}$ from DGM $p_{\theta_{pt}}$ are used as initial weights $\theta_{0}$ to train the generative model $p_{\hat{\theta}}$ (see Fig. 1).
Figure 2: Block diagram for the pre-training case. The inductive bias is introduced by training a DGM $p_{\theta_{pt}}$ on a large collection of $x_{g}$ samples. The weights learned from this training process with abundant samples serve as the initial parameters $\hat{\theta}$ for the fine-tuning process using the real data $x_{r}$ to obtain $p_{\hat{\theta}}$.
In essence, our approach aligns with the well-established concept of data augmentation. We generate synthetic data $x_{g}$, which may not perfectly capture the intricacies of the original data $x_{r}$. However, we used these synthetic data to train another DGM $p_{\theta_{pt}}$. Although $p_{\theta_{pt}}$ might generate lower-quality synthetic samples, our objective is to exploit the information encoded within this DGM to establish an initial set of weights for the DGM that will eventually be trained on $x_{r}$. In other words, we exploit the knowledge of the generative model $p_{\theta_{pt}}$, the context domain, to obtain a better generative model $p_{\hat{\theta}}$, which is our target domain. Fig. 2 visually represents this pre-training procedure.
@@ 2.1.2 Model Averaging
The concept of model averaging emerged in the 1960s, primarily within the field of economics (model_averaging, ; d890a735-c271-388b-8b71-a309e52af62b, ). Traditional empirical research often selects a single “best" model after searching a wide space of possibilities. However, this approach can underestimate the real uncertainty, leading to overly confident conclusions. Model averaging offers a compelling alternative. By combining multiple models, the resulting ensemble can outperform any individual model. This approach aligns with the core principles of statistical modeling, which involve maximizing information use while balancing flexibility with the risk of overfitting. In essence, model averaging extends the concept of model selection by leveraging insights from all the models considered.
While pre-training can be incorporated with any DGM, our approach focuses on models where the training process is sensitive to initial conditions, such as VAEs. In such cases, it is common to train the DGM $p_{\theta}$ with multiple initial conditions (seeds) and potentially discard “bad" seeds based on a specific metric. We propose using these discarded seeds to create an artificial inductive bias. The simplest implementation involves averaging the model parameters. In this case, our context domains are the different results obtained from each seed, and the target domain is obtained by averaging across these context domains. If we train $S$ different seeds for $p_{\theta}$, resulting in $S$ models with parameters $\theta_{s}$, we propose using the average of these weights as the inductive bias:
@@
This straightforward approach is computationally efficient, requiring only the calculation of the average across the precomputed weights. It assumes that the average model may capture a robust inductive bias, leading to improved performance. Fig. 3 summarizes this process.
Figure 3: Block diagram for the model averaging case. The inductive bias is introduced by training a DGM $p_{\theta}$ on $x_{r}$ using $S$ different seeds. The average of the weights learned from these training processes serves as the initial parameters $\hat{\theta}$ for the fine-tuning process using the real data $x_{r}$ to obtain $p_{\hat{\theta}}$.
@@ 2.2 Meta-learning
Traditional machine learning models often rely on large volumes of data to achieve optimal performance in specific tasks. In contrast, meta-learning introduces a distinct paradigm by training algorithms that can “learn to learn" (Thrun1998, ), enabling them to rapidly adapt to new tasks with minimal data. This departure from the conventional requirement of extensive datasets for each new task allows meta-learning algorithms to leverage knowledge by addressing numerous related tasks. Through introspective analysis of past experiences, these models dynamically adjust their learning strategies when confronted with novel situations, making them more efficient learners and requiring less data to perform well on tasks with similar characteristics.
In this work, we exploit the multi-seed training configuration of certain DGMs. We construct a meta-learning framework by treating each $S$ different seeds obtained after training the DGM as a distinct task.
@@ 2.2.1 MAML
MAML is a prevalent approach within the field of meta-learning (finn2017model, ). It identifies the initial set of weights denoted by $\theta_{MAML}$ by leveraging various tasks, enabling rapid and data-efficient adaptation to new tasks. This efficiency comes from fine-tuning the $\theta_{MAML}$ with minimal data for each new task. However, the successful application of MAML requires access to diverse tasks for effective learning.
Formally, we can frame the problem by starting with a common single-task learning scenario and transforming it into a meta-learning framework. Consider a task $\mathcal{T}$ that consists of an input $x$ sampled from a probability distribution $\mathcal{D}$. For simplicity, we define a task instance $\mathcal{T}$ as a tuple comprising a dataset $\mathcal{D}$ and its corresponding loss function $\mathcal{L}$. To solve the task $\mathcal{T}$, we need to obtain an optimal model parameterized by a task-specific parameter $\omega^{*}$, which minimizes a loss function $\mathcal{L}$ on the data of the task as follows:
In single-task learning, hyperparameter optimization is achieved by splitting the dataset $\mathcal{D}$ into two disjoint subsets $\mathcal{D}=\mathcal{D}^{(t)}\cup\mathcal{D}^{(v)}$, which are the training and validation sets, respectively. The meta-learning setting aims to develop a general-purpose learning algorithm that excels across a distribution of tasks represented by $p(\mathcal{T})$ (Hospedales2020MetaLearningIN, ). The objective is to use training tasks to train a meta-learning model $\theta_{MAML}$ that can be fine-tuned to obtain $\omega$ to perform well on unseen tasks sampled from the same task environment $p(\mathcal{T})$. Meta-learning methods utilize meta-parameters to model the common latent structure of the task distribution $p(\mathcal{T})$. Therefore, we consider meta-learning an extension of hyperparameter optimization, where the hyperparameter of interest – often referred to as a meta-parameter – is shared across multiple tasks.
In this work, the distribution of tasks is defined by the set of $S$ training seeds obtained after training the DGM. Given a set of $S$ training seeds following $p(\mathcal{T})$, each task $\mathcal{T}\sim p(\mathcal{T})$ is therefore formalized as $\mathcal{T}=\{\mathcal{D},\mathcal{L}\}$. Each dataset $\mathcal{D}$ consists of synthetic data points $x_{g}^{s}$ drawn from the model for the different training seeds. The loss function $\mathcal{L}$ corresponds to the DGM loss function. The specific $\mathcal{L}$ form depends on the chosen DGM. If the chosen DGM is a VAE, the loss function $\mathcal{L}$ would be the negative of the Evidence Lower BOund (ELBO) (Kingma2013AutoEncodingVB, ). In contrast, if a GAN is used, the loss function $\mathcal{L}$ would be the minimax loss function arising from the interplay between the generator and discriminator networks (Goodfellow2014GenerativeAN, ). It is essential to note that both VAEs and GANs employ two neural networks within their architecture, which differs from the single network architectures commonly found in state-of-the-art applications (Gordon2018MetaLearningPI, ; SmithMiles2009CrossdisciplinaryPO, ).
@@
Solving this problem using the MAML approach requires access to a collection of $B$ tasks sampled from $p(\mathcal{T})$. We denote this set of tasks $\mathcal{T}_{b}$ used for training as $\mathcal{D}_{b}=\{(\mathcal{D}_{b}^{(t)},\mathcal{D}^{(v)}_{b})\}_{b=1}^{B}$, where each task $b$ has dedicated meta-training and meta-validation data, respectively. The goal of meta-training is to find the optimal $\omega^{*}_{b}$ for a given task $b$ given $\theta_{MAML}$. This $\theta_{MAML}$ essentially captures the ability to learn effectively from new data. In this context, the task-related parameter $\omega_{b}$ refers to the task-specific parameters of the two networks that comprise the VAE, namely, the encoder and decoder. After meta-training, the learned $\omega^{*}_{b}$ is used to guide the training of a base model $\theta_{MAML}$. This procedure is called meta-testing. This essentially means that the model leverages the knowledge gained from previous tasks to improve the efficiency of learning on new tasks. This can be viewed as a bi-level optimization problem (Franceschi2018BilevelPF, ):
@@
This equation minimizes the expected loss across all tasks on the meta-validation sets, subject to the constraint that the task-specific parameter $\omega$ is optimized on the corresponding meta-training data for each task.
Since in our work, we are upgrading the parameters using gradient descent, we can reformulate Equation˜4 as follows:
@@
Fig. 4 illustrates the integration of the MAML procedure within the framework of our proposed methodology. In this context, the task space denoted by $p(\mathcal{T})$ corresponds to the various seeds $S$ obtained during the training process. Essentially, the task space encompasses the different probability distributions $p_{\theta_{s}}$ associated with each training seed. Ultimately, the meta-training steps lead to identifying the desired parameters, denoted by $\theta_{MAML}$. Note that $\theta_{MAML}$ represents a set of parameters that adapt fast to new data; in our case, it means that the DGM initial parameters are chosen so that they adapt fast to generate real data.